A force on a particle due to electric and magnetic field is given by
F=qE+q(v×B).
Suppose E is along the x-axis and B is along y-axis. In what direction and at what minimum speed v should a positively charged particle be sent so that the net force on it is zero? Plz i dont get this question .plz help me 

Grade:12th pass

2 Answers

Nicho priyatham
8 years ago
along +ve z axis
v=E/B
PLZ APPROVE if useful and incase of anydoubts put it is ans box
Saiman Qureshi
8 years ago
How v=e/b
Acc to ques F=0
ie qE+q(v*B)=0
E=-(v*B)
E=-vbsin@
v=-E/b.sin@